OpenStudy (anonymous):

u3 - v3 Can you show me how to do this?

OpenStudy (anonymous):

u should know that : \[a^3-b^3=(a-b)(a^2+ab+b^2)\] put u and v at a and b respectively \[u^3-v^3=(u-v)(u^2+uv+v^2)\]